Early this week (7- 9 June), GONG/LE observations were interrupted by a damage to the camera trigger cable. The root-cause was identified by the NISP engineering team via remote troubleshooting, and while the cable cannot be repaired, the solution was found to restore the operations. During the troubleshooting, the H-alpha observations may show a rapid image rotation. The observations are back to normal starting 11 June, 2025.
Subscribe to:
Post Comments (Atom)
X-ray flare in a growing active region
Active region 14366 crossed East solar limb at about 18:00 UT on 29 January, 2026. The region was growing rapidly, and it was a location of ...
-
The peakfind results for GONG Month 302 show that the frequency shifts have started to decrease indicating that the solar maximum has alrea...
-
GONG/Learmonth (Australia) took observations of a large two-ribbon flare, which was the source of Earth-directed halo Coronal Mass Ejectio...
-
A notification was posted on the BVES website regarding the outage: Southern California Edison (SCE)- Restoration and Continued Public S...
No comments:
Post a Comment